OpenStudy (anonymous):

what is the derivative of f(x)=2sqrt(x+(1/4))?

OpenStudy (abb0t):

\[2\sqrt{x+\frac{1}{4}} = 2 \frac{d}{dx} [x+\frac{1}{4}]^{\frac{1}{2}}\] does that help? You'll be using power and chain rule.

OpenStudy (anonymous):

so it would be \[f'(x)=\frac{ 1 }{ x+\frac{ 1 }{ 4 } }\] ?

OpenStudy (anonymous):

shoot that's not what i meant. i meant for the x+1/4 to be under a square root

OpenStudy (abb0t):

yes.

OpenStudy (anonymous):

awesome, thank you so much!